# Mapping the Cloud-Driven Atmospheric Dynamics & Chemistry of an Isolated Exoplanet Analog with Harmonic Signatures

### v3

@@ 4.1 Harmonic Analysis
To investigate the rotational modulation and atmospheric structure of SIMP J0136, we applied 1-D harmonic wave decomposition (see §3.1) to the lightcurves obtained from both the NIRISS/SOSS and NIRSpec/PRISM instruments. By decomposing the observed variability into sinusoidal components, we retrieved wave periodicity, amplitude, phase shift, and damping ratios as a function of wavelength and associated atmospheric depth. These results allow us to examine how different spectral regions and molecular absorption (and potentially emission) features contribute to the observed spectrophotometric variability of SIMP J0136. By testing models with varying harmonic complexity, we assess the presence of layered atmospheric dynamics and chemical structures.
4.1.1 NIRISS/SOSS Harmonic Results
@@
Considering NIRISS/SOSS lightcurves that span from 1.0 to 2.8 m with 0.05 m intervals (36 lightcurves), we identify periodic signals for each bin by fitting harmonic models using the Bayesian framework described in §3. We assume that variations in specific wavelength intervals corresponding to known molecular bands reflect variations in the absorption of those species. If the observed molecular variability was only due to the presence of clouds muting spectral features, all species would be uniformly affected. However, we see in §4.2.2 that the presence of clouds leads to enhanced absorption for methane and reduced absorption for H2O and CO. This allows us to connect the retrieved rotational periodicities, amplitudes, and phase shifts with changes in atmospheric composition and depth.
@@
The lowest-order periodicity across NIRISS/SOSS wavelengths is the harmonic at 2.4 h, consistent with SIMP J0136’s rotation (Yang2016). These harmonics tend to have the highest uncertainty as the period is of similar order as the observational period.
CO bands, and to a lesser extent H2O, show retrieved periods that consistently fall between the fundamental rotation (, 2.4 h) and its first harmonic (, 1.2 h), producing beating patterns with characteristic periodicities spanning 1.2 to 2.4 h at pressure levels ranging from to bar. These beating patterns were also seen in Apai2017; apai21; Fuda:2024 for WISE 1049 B and Zhou2022 for VHS 1256 b.
Higher-order () variability traces deeper atmospheric layers as seen in Figure 3. The pressure levels associated with the maximum flux from the spectral contribution function (see §3.4) are overlaid onto the retrieved components. From 1.0 to 1.8 m and to 2.2 m, the (0.8 h) harmonic is preferentially retrieved at pressure levels bar, consistent with variability originating from deeper atmospheric layers. It is noteworthy that the harmonics also have the smallest posterior uncertainty.
@@
4.1.2 NIRSpec/PRISM Harmonic Results
@@
4.2 Vertical Variability Mapping via Harmonic Waves
Time-resolved, multi-spectral data such as those collected by JWST NIRISS and NIRSpec instruments provide the opportunity to map the dynamic atmospheric physics of brown dwarfs and exoplanets at varying pressure levels. Using a method first demonstrated in Akhmetshyn2025, we use the best-fit harmonic models (Figure 2, Middle Row), along with a model contribution function (see §3.4), to create vertical flux variability heat maps (pressure vs. time) of the atmosphere of SIMP J0136. These vertical variability maps demonstrate the flux variability at each atmospheric layer, the interaction between layers, and unique spectral contributions to the object’s dynamic atmosphere.
4.2.1 Composite Vertical Variability Maps
To build composite vertical variability maps, we start by scaling the model contribution function, , see Equation 5, by the normalized best-fit lightcurve model, , see Equation 1, for each spectral bin, , and time-step. At each pressure level, , we then numerically sum the scaled spectral contribution across wavelength bins to compute the pressure- and time- dependent variability,
resulting in a 1-D flux array in pressure-level space. The computation is then repeated at each time-step, followed by normalization of flux at each pressure level, resulting in a composite vertical variability map (Figure 5, Top Row) in terms of % deviation. The composite deviation maps broadly match the dynamic spectra (variability maps) in Figure 2, and the vertical variability maps can be seen as a coordinate transformation of the flux heat map from time vs. wavelength to time vs. pressure.
@@ 4.2.1 Composite Vertical Variability Maps
SIMP J0136’s composite vertical variability maps of the NIRISS/SOSS and NIRSpec/PRISM data reveal both similarities in the atmospheric structure observed at each epoch and apparent qualitative differences between the two observations (separated by h or rotations). Both NIRISS/SOSS and NIRSpec/PRISM maps show two dynamic interacting atmospheric layers with time-varying boundaries: (1) a lower pressure (higher altitude) overlying layer comprised of lower order harmonics and (2) a deeper layer (bar) corresponding to expected cloud formation regions of forsterite and potentially iron (see e.g., Vos2023; McCarthy:2024; McCarthy2025). The time-varying pressure level boundary between the deep and overlying atmospheric layers suggests variations in vertical cloud structure.
@@
4.2.2 Spectral Features’ Deviations from Bulk Variability
The broad NIR spectral coverage of JWST NIRISS/SOSS and NIRSpec/PRISM allows us to map spectral features’ deviations from bulk (or composite) variability and investigate that deviation’s ties to distinct atmospheric phenomena. For example, we can determine how wavelengths corresponding to iron ( to m) and forsterite ( to m) cloud modulation (Vos2023; McCarthy2025) or known molecular absorption bands (H2O, CO, CH4) deviate from bulk variability over the period of observation.
@@ 4.2.2 Spectral Features’ Deviations from Bulk Variability
It is important to note that maps showing deviations of individual spectral features from the bulk variability do not necessarily indicate bona fide, feature-specific variability. Rather, they reflect differences in the flux relative to the composite (bulk) signal. For example, a molecular absorption band may exhibit no intrinsic time-dependent behavior, but if bulk variability is subtracted from the spectral band light curves, the residuals would exhibit an induced anti-correlation. Nevertheless, these deviations from the bulk variability are indeed meaningful, as they highlight wavelengths with distinct morphologies from the overall spectrum.
The deviation maps for wavelengths expected to correspond to mineral cloud modulation (i.e., iron and forsterite, Vos2023) demonstrate clear harmonic patterns for both the NIRISS/SOSS and NIRSpec/PRISM datasets (see Figure 6, Top Row). The lower layer (bar) is suspected to be iron cloud modulation with maximum signal-to-noise (S/Nmax) deviations of 18.5 for NIRISS/SOSS and 16.1 for NIRSpec/PRISM observations. The upper forsterite layer (centered at 1 bar) is approximately anti-correlated with the underlying iron layer and has S/Nmax deviations of 16.3 and 26.5 for NIRISS/SOSS and NIRSpec/PRISM respectively.
For the forsterite variation, the NIRISS/SOSS data include a double-peaked dark patch but the NIRSpec/PRISM observations contain only a single dark patch. The iron cloud regions do not appear to possess double-peaked features. For these maps, the dark and bright patches are presumed to be longitudes of enhanced or reduced cloud coverage, respectively, and will be discussed further in §5.2.
@@
5 Discussion
5.1 Odd Harmonics in the Forsterite Cloud Layer
@@
As initially reported by Akhmetshyn2025, odd () harmonics in the NIRISS/SOSS data may indicate North/South asymmetry in the surface brightness map of SIMP J0136 (depending on the limb darkening law; luger21d). In the absence of limb darkening, even spherical-harmonic degrees () paired with odd azimuthal orders (which correspond to the longitudinal wavenumber ) signify a North–South asymmetry in the surface-brightness distribution (Cowan&Fuentes2013), an assumption adopted by Akhmetshyn2025. The situation changes, however, once limb darkening is included: linear and quadratic limb–darkening laws effectively increase the contributing spherical-harmonic degree by one and two, respectively (luger21d). Consequently, a linear limb–darkening law can generate apparent odd- (odd-) signals even for a North–South symmetric map, whereas under a quadratic law—or in the limit of negligible limb darkening—the presence of such odd modes once again implies genuine hemispheric asymmetry. Additional constraints on the limb–darkening characteristics of SIMP J0136 and similar objects will therefore be essential for robust interpretation of their light-curve morphologies.
If the odd harmonics are indeed due to North/South asymmetry, then they also support an inclined viewing geometry (Cowan&Fuentes2013), consistent with the estimate of (Vos2017). Furthermore, this asymmetry could help lift hemispheric ambiguity in future Doppler imaging efforts with ELT-class telescopes with high-resolution spectrographs (see e.g., Plummer&Wang2023), making SIMP J0136 an ever more compelling target for such work.
Because brown dwarfs have been found to be redder, and therefore cloudier, at equatorial latitudes than polar ones (Vos2017; Suarez&Metchev2022; Suarez2023), the odd harmonics at the deepest pressure levels (corresponding to the cloud formation region) potentially suggest the North/South asymmetry may lie in the mineral cloud modulation region and be concentrated near the equator. Multi-rotational NIR spectroscopic observations of SIMP J0136 would help to constrain higher order odd harmonics, decrease the uncertainty in retrieved harmonic parameters, and elucidate the dynamical structure of L/T transition objects.
5.2 Cloud-Driven Temperature and Chemical Structure
@@
Forsterite cloud modulation can be seen to drive the upper atmosphere temperature structure as shown by the correlation between the dark cloud feature in Figure 6 (Top Row, Right Column) and the high-altitude continuum in Figure 6 (Bottom Row, Right Column). Regions above the clouds are relatively cooler while gaps in the cloud layer allow heat to escape initially in vertical columns (see e.g., Showman&Tan2019; tan21a; tan21b). As CO is the preferred carbon reservoir at higher temperatures (and contrastingly CH4 is the preferred reservoir at lower temperatures), cooler regions should exhibit depleted CO and enhanced CH4 absorption, matching our findings (Figure 6, Third and Fourth Rows) and contributing to carbon disequilibrium chemistry in SIMP J0136.
To understand the H2O deviations in this context, we should consider the following equation (e.g., Visscher2011),
@@
The sequestration of oxygen in the forsterite cloud formation process may also factor into the observed anti-correlation of the cloud layer and CO/H2O. The chemistry governing the forsterite cloud formation process can be understood by the net thermochemical reaction (e.g., Visscher2010),
@@
5.3 High-Altitude Heating Source
@@
Mapping the deviations in the wavelengths (3.05 to 3.5 m) bracketing the high-altitude, rotational-vibrational CH4 band (m, see Figure 6, Fifth Row) results in harmonics distinct from and largely anti-correlated with the low-altitude CH4 bands (see Figure 6, Fourth Row). For example, the CH4 absorption feature between 38.0 and 38.5 hr on the low-altitude CH4 map is replaced by a bright spot, which we tentatively interpret as an emission feature in the upper atmosphere, potentially driven by auroral heating from electron precipitation (e.g., Sandel1982; Melin2011; Muller-Wodarg2012; Pineda2024; McCarthy2025; Nasedkin2025). In this case, the CH4 abundance appears variable and is influenced and in-phase with underlying cloud modulation, but now we detect the CH4 in emission rather than absorption.
@@
Distinct, high-altitude CH4 emission also aligns with previous findings for SIMP J0136 based on the same NIRSpec/PRISM dataset explored in this work. K-means clustering identified unique lightcurve morphology near the CH4 -band (McCarthy2025), and a 265 K temperature inversion was inferred at pressures (bar) corresponding to the same wavelengths (Nasedkin2025). The necessary power to create such a strong temperature inversion from electron precipitation-driven auroral heating (which is similar in strength to the 300 K heating found for WISE J1935, Faherty2024), was computed to be W by Nasedkin2025 while Jupiter-like auroral processes could only produce auroral power on the scale of W.
@@
As discussed in Nasedkin2025, other upper atmospheric heating processes such as the breaking of gravity waves (Freytag2010) and Joule heating associated with atmospheric electric currents have been proposed for brown dwarfs based on observations of these mechanisms in outer solar system planets (Cowley2004; Ingersoll2021; Brown2022). Further studies will be necessary to break the degeneracies between these potential heating sources.
6 Summary

### v4

@@ IV.1 Harmonic Analysis
To investigate the rotational modulation and atmospheric structure of SIMP J0136, we applied 1-D harmonic wave decomposition (see §III.1) to the lightcurves obtained from both the NIRISS/SOSS and NIRSpec/PRISM instruments. By decomposing the observed variability into sinusoidal components, we retrieved wave periodicity, amplitude, phase shift, and damping ratios as a function of wavelength and associated atmospheric depth. These results allow us to examine how different spectral regions and molecular absorption (and potentially emission) features contribute to the observed spectrophotometric variability of SIMP J0136. By testing models with varying harmonic complexity, we assess the presence of layered atmospheric dynamics and chemical structures.
IV.1.1 NIRISS/SOSS Harmonic Results
@@
Considering NIRISS/SOSS lightcurves that span from 1.0 to 2.8 m with 0.05 m intervals (36 lightcurves), we identify periodic signals for each bin by fitting harmonic models using the Bayesian framework described in §III. We assume that variations in specific wavelength intervals corresponding to known molecular bands reflect variations in the absorption of those species. If the observed molecular variability was only due to the presence of clouds muting spectral features, all species would be uniformly affected. However, we see in §IV.2.2 that the presence of clouds leads to enhanced absorption for methane and reduced absorption for H2O and CO. This allows us to connect the retrieved rotational periodicities, amplitudes, and phase shifts with changes in atmospheric composition and depth.
@@
The lowest-order periodicity across NIRISS/SOSS wavelengths is the harmonic at 2.4 h, consistent with SIMP J0136’s rotation (Yang et al., 2016). These harmonics tend to have the highest uncertainty as the period is of similar order as the observational period.
CO bands, and to a lesser extent H2O, show retrieved periods that consistently fall between the fundamental rotation (, 2.4 h) and its first harmonic (, 1.2 h), producing beating patterns with characteristic periodicities spanning 1.2 to 2.4 h at pressure levels ranging from to bar. These beating patterns were also seen in Apai et al. (2017, 2021b); Fuda et al. (2024) for WISE 1049 B and Zhou et al. (2022) for VHS 1256 b.
Higher-order () variability traces deeper atmospheric layers as seen in Figure 3. The pressure levels associated with the maximum flux from the spectral contribution function (see §III.4) are overlaid onto the retrieved components. From 1.0 to 1.8 m and to 2.2 m, the (0.8 h) harmonic is preferentially retrieved at pressure levels bar, consistent with variability originating from deeper atmospheric layers. It is noteworthy that the harmonics also have the smallest posterior uncertainty.
@@
IV.1.2 NIRSpec/PRISM Harmonic Results
@@
IV.2 Vertical Variability Mapping via Harmonic Waves
Time-resolved, multi-spectral data such as those collected by JWST NIRISS and NIRSpec instruments provide the opportunity to map the dynamic atmospheric physics of brown dwarfs and exoplanets at varying pressure levels. Using a method first demonstrated in Akhmetshyn et al. (2025), we use the best-fit harmonic models (Figure 2, Middle Row), along with a model contribution function (see §III.4), to create vertical flux variability heat maps (pressure vs. time) of the atmosphere of SIMP J0136. These vertical variability maps demonstrate the flux variability at each atmospheric layer, the interaction between layers, and unique spectral contributions to the object’s dynamic atmosphere.
IV.2.1 Composite Vertical Variability Maps
To build composite vertical variability maps, we start by considering a time-varying contribution function of similar form as Equation 5,
Adopting a similar approach as that seen in phase-curve mapping literature (e.g., Irwin et al., 2020; Al-Refaie et al., 2021; Changeat et al., 2021; Changeat, 2022; Chubb and Min, 2022), we assume that opacity () is time-independent and that temporal or longitudinal variability is only expressed through the source function (a necessary assumption so that we can employ forward, grid-based, and self-consistent atmospheric models) such that
and therefore,
resulting in a pressure- and wavelength-dependent transmission term,
We can then restate Equation 6 as follows,
Again borrowing from phase-curve mapping literature, we approximate the source function as a Planck function, , and linearize about a reference temperature, (see e.g., Mihalas and Mihalas, 1984; Goody and Yung, 1989). We then parameterize the components of the temperature perturbation, , as separable variables (see e.g., Rodgers, 2000). Linearization requires that , a requirement met by the small % variability in our data. The result of this parameterization is that the source function is correspondingly factorized as follows:
where represents a component of the mean atmospheric temperature structure and , the spectral variability of the emitted radiation. The spectral flux, , is modeled using , the normalized, best-fit spectral lightcurves (see Equation 1).
Combining Equations 10 and 11 results in a time-varying contribution function in which the time-dependence is isolated to the spectral flux, ,
and it can be seen that is the time-independent (mean) contribution function, .
With these results in mind, we compute the pressure- and time-dependent variability, , at each time-step and pressure level () and numerically sum the scaled spectral contribution across wavelength bins,
resulting in a 1-D flux array in pressure-level space. The computation is then repeated at each time-step, followed by normalization of flux at each pressure level, resulting in a composite vertical variability map (Figure 5, Top Row) in terms of % deviation. The composite deviation maps broadly match the dynamic spectra (variability maps) in Figure 2, and the vertical variability maps can be seen as a (model-weighted) coordinate transformation of the flux heat map from time vs. wavelength to time vs. pressure.
@@ IV.2.1 Composite Vertical Variability Maps
SIMP J0136’s composite vertical variability maps of the NIRISS/SOSS and NIRSpec/PRISM data reveal both similarities in the atmospheric structure observed at each epoch and apparent qualitative differences between the two observations (separated by h or rotations). Both NIRISS/SOSS and NIRSpec/PRISM maps show two dynamic interacting atmospheric layers with time-varying boundaries: (1) a lower pressure (higher altitude) overlying layer comprised of lower order harmonics and (2) a deeper layer (bar) corresponding to expected cloud formation regions of forsterite and potentially iron (see e.g., Vos et al., 2023; McCarthy et al., 2024, 2025). The time-varying pressure level boundary between the deep and overlying atmospheric layers suggests variations in vertical cloud structure.
@@
IV.2.2 Spectral Features’ Deviations from Bulk Variability
The broad NIR spectral coverage of JWST NIRISS/SOSS and NIRSpec/PRISM allows us to map spectral features’ deviations from bulk (or composite) variability and investigate that deviation’s ties to distinct atmospheric phenomena. For example, we can determine how wavelengths corresponding to iron ( to m) and forsterite ( to m) cloud modulation (Vos et al., 2023; McCarthy et al., 2025) or known molecular absorption bands (H2O, CO, CH4) deviate from bulk variability over the period of observation.
@@ IV.2.2 Spectral Features’ Deviations from Bulk Variability
It is important to note that maps showing deviations of individual spectral features from the bulk variability do not necessarily indicate genuine, feature-specific variability. They reflect differences in the flux relative to the composite (bulk) signal. For example, a molecular absorption band may exhibit no intrinsic time-dependent behavior, but if bulk variability is subtracted from the associated dynamic spectra, the resulting residuals would exhibit an induced anti-correlation. Despite this, deviations from the bulk variability can be meaningful and provide insight into wavelengths with distinct time-varying behavior.
The deviation maps for wavelengths expected to correspond to mineral cloud modulation (i.e., iron and forsterite, Vos et al., 2023) demonstrate clear harmonic patterns for both the NIRISS/SOSS and NIRSpec/PRISM datasets (see Figure 6, Top Row). The lower layer (bar) is suspected to be iron cloud modulation with maximum signal-to-noise (S/Nmax) deviations of 18.5 for NIRISS/SOSS and 16.1 for NIRSpec/PRISM observations. The upper forsterite layer (centered at 1 bar) is approximately anti-correlated with the underlying iron layer and has S/Nmax deviations of 16.3 and 26.5 for NIRISS/SOSS and NIRSpec/PRISM respectively.
For the forsterite variation, the NIRISS/SOSS data include a double-peaked dark patch but the NIRSpec/PRISM observations contain only a single dark patch. The iron cloud regions do not appear to possess double-peaked features. For these maps, the dark and bright patches are presumed to be longitudes of enhanced or reduced cloud coverage, respectively, and will be discussed further in §V.2.
@@
V Discussion
V.1 Odd Harmonics in the Forsterite Cloud Layer
@@
As initially reported by Akhmetshyn et al. (2025), odd () harmonics in the NIRISS/SOSS data may indicate North/South asymmetry in the surface brightness map of SIMP J0136 (depending on the limb darkening law; Luger et al., 2021). In the absence of limb darkening, even spherical-harmonic degrees () paired with odd azimuthal orders (which correspond to the longitudinal wavenumber ) signify a North–South asymmetry in the surface-brightness distribution (Cowan et al., 2013), an assumption adopted by Akhmetshyn et al. (2025). The situation changes, however, once limb darkening is included: linear and quadratic limb–darkening laws effectively increase the contributing spherical-harmonic degree by one and two, respectively (Luger et al., 2021). Consequently, a linear limb–darkening law can generate apparent odd- (odd-) signals even for a North–South symmetric map, whereas under a quadratic law—or in the limit of negligible limb darkening—the presence of such odd modes once again implies genuine hemispheric asymmetry. Additional constraints on the limb–darkening characteristics of SIMP J0136 and similar objects will therefore be essential for robust interpretation of their light-curve morphologies.
If the odd harmonics are indeed due to North/South asymmetry, then they also support an inclined viewing geometry (Cowan et al., 2013), consistent with the estimate of (Vos et al., 2017). Furthermore, this asymmetry could help lift hemispheric ambiguity in future Doppler imaging efforts with ELT-class telescopes with high-resolution spectrographs (see e.g., Plummer and Wang, 2023), making SIMP J0136 an ever more compelling target for such work.
Because brown dwarfs have been found to be redder, and therefore cloudier, at equatorial latitudes than polar ones (Vos et al., 2017; Suárez and Metchev, 2022; Suárez et al., 2023), the odd harmonics at the deepest pressure levels (corresponding to the cloud formation region) potentially suggest the North/South asymmetry may lie in the mineral cloud modulation region and be concentrated near the equator. Multi-rotational NIR spectroscopic observations of SIMP J0136 would help to constrain higher order odd harmonics, decrease the uncertainty in retrieved harmonic parameters, and elucidate the dynamical structure of L/T transition objects.
V.2 Cloud-Driven Temperature and Chemical Structure
@@
Forsterite cloud modulation can be seen to drive the upper atmosphere temperature structure as shown by the correlation between the dark cloud feature in Figure 6 (Top Row, Right Column) and the high-altitude continuum in Figure 6 (Bottom Row, Right Column). Regions above the clouds are relatively cooler while gaps in the cloud layer allow heat to escape initially in vertical columns (see e.g., Showman et al., 2019; Tan and Showman, 2021a, b). As CO is the preferred carbon reservoir at higher temperatures (and contrastingly CH4 is the preferred reservoir at lower temperatures), cooler regions should exhibit depleted CO and enhanced CH4 absorption, matching our findings (Figure 6, Third and Fourth Rows) and contributing to carbon disequilibrium chemistry in SIMP J0136.
To understand the H2O deviations in this context, we should consider the following equation (e.g., Visscher and Moses, 2011),
@@
The sequestration of oxygen in the forsterite cloud formation process may also factor into the observed anti-correlation of the cloud layer and CO/H2O. The chemistry governing the forsterite cloud formation process can be understood by the net thermochemical reaction (e.g., Visscher et al., 2010),
@@
V.3 High-Altitude Heating Source
@@
Mapping the deviations in the wavelengths (3.05 to 3.5 m) bracketing the high-altitude, rotational-vibrational CH4 band (m, see Figure 6, Fifth Row) results in harmonics distinct from and largely anti-correlated with the low-altitude CH4 bands (see Figure 6, Fourth Row). For example, the CH4 absorption feature between 38.0 and 38.5 hr on the low-altitude CH4 map is replaced by a bright spot, which we tentatively interpret as an emission feature in the upper atmosphere, potentially driven by auroral heating from electron precipitation (e.g., Sandel et al., 1982; Melin et al., 2011; Müller-Wodarg et al., 2012; Pineda et al., 2024; McCarthy et al., 2025; Nasedkin et al., 2025). In this case, the CH4 abundance appears variable and is influenced and in-phase with underlying cloud modulation, but now we detect the CH4 in emission rather than absorption.
@@
Distinct, high-altitude CH4 emission also aligns with previous findings for SIMP J0136 based on the same NIRSpec/PRISM dataset explored in this work. K-means clustering identified unique lightcurve morphology near the CH4 -band (McCarthy et al., 2025), and a 265 K temperature inversion was inferred at pressures (bar) corresponding to the same wavelengths (Nasedkin et al., 2025). The necessary power to create such a strong temperature inversion from electron precipitation-driven auroral heating (which is similar in strength to the 300 K heating found for WISE J1935, Faherty et al., 2024), was computed to be W by Nasedkin et al. (2025) while Jupiter-like auroral processes could only produce auroral power on the scale of W.
@@
As discussed in Nasedkin et al. (2025), other upper atmospheric heating processes such as the breaking of gravity waves (Freytag et al., 2010) and Joule heating associated with atmospheric electric currents have been proposed for brown dwarfs based on observations of these mechanisms in outer solar system planets (Cowley et al., 2004; Ingersoll et al., 2021; Brown et al., 2022). Further studies will be necessary to break the degeneracies between these potential heating sources.
VI Summary